AC Circuit Analysis3. Three balanced three-phase loads are connected in parallel. Load 1 is Y- connected with an impedance of 400 + j300N/0 load
Answer #1

ANS:

First calculate active a reactive power of all the loads.

(1) Y-connected load

Load impedance per phase Zy 400+ 7300

  300 = 4002 + 3002 Ztan 400

  500236,862

Line to neutral voltage at the load end  V = 16V3kV

Current in the phase of Y-connected load,

P Іү Zy

  16V3 x 103 500

55.42A

Active power in the load per phase = 1} Ry

  = 55.422 x 400

= 1228.55k

Total active power of load Py = 3 x 1228.55 = 3685.65kW   

Reactive power in the load per phase = 1} Xy

  = 55.424 х 300

921.41KV Ar

Total reactive power in the load Qy = 3 x 921.41 = 2764.33kV Ar

Complex power of the Y connected load ,

Sy Py +jQy

Sy = 3685.65 + 12764.33kVA

(2)

Delta connected load

Delta connected load per phase - 2400 - 1800Ω

  = 24002 + 180022 – tan 1800 2400

  სი989$ -7ბბბ¢ =

Line to neutral voltage at the load end  V = 16V3kV

Load is balanced, therefore line voltage VL = 3V P

= V3 x 16V3

= 48kV

Magnitude of phase current in delta connected load

IA VL ZA

  48 x 103 3000

= 16A

Active power in the load per phase 1² RA

  = 162 х 2400

= 61ᏎᏎᎬᏉ

Total active power of load PA = 3 X 614.4= 1843.2kW   

Reactive power in the load per phase = 1 X4

  = 162 x (-1800

-460.8KV Ar

Total reactive power in the load QA = 3 X (-460.8) = -1382.4kV Ar

Complex power of the delta connected load ,

SA PA+QA

SA = 1843.2 - 31382.4kV A

(3)

Load 3 S3 = 270 +79000KV A

Pa = 270kW   Q3 9000KV Ar

Total active power of all load Pload Py + PA+ P3

=3685.65+1843.2+270

  =5798.85kW

Total reactive power of all loads Q_{load}=Q_Y+Q_{\Delta}+Q_3

=2764.33-1382.4+9000

=10381.93kVAr

Complex powe of load Sload Pload + iQload

= 5798.85 +j10381,Ꮽ3ᏂᏙᎪ

Reactive Power (Q) Apparant Power (S) Active Power (P)​​​​​​​

Powefactor angle of combined load ,

\phi=tan^{-1}\frac{Q_{load}}{P_{load}}

\phi=tan^{-1}\frac{10381.93}{5798.85}

  = 60.81

Recall the equation of power for three phase balanced load

P=\sqrt3V_LI_Lcos(\phi)

(o)soɔTAE IL = P

Put P=P_{load}=5798.85kW   V_L=48kV  \phi=60.81^0

I_L=\frac{P}{\sqrt3V_Lcos(\phi)}

=\frac{5798.85\times10^3}{\sqrt3\times48\times10^3cos(60.81^0)}

= 143.01A

Now find complex power of line .

Line impedance per phaseZ_L=1+j4\Omega

S_{line}=3I_L^2Z_L​​​​​​​

  =3\times 143.01^2 (1+j4)

  =61.355+j245.422kVA

(a)

Total complex power at the source,

Ss = SOUT ce Sline + Sload

  =61.355+j245.422+5798.85+j10381.93

  5860.205+310627352EVA

(b)

Average power refers to active power.

Active power at source Ps = 5860.205kW

Active power at load P_{load}=5798.85kW

%Average power delivered to load =  \frac{P_{load}}{P_S}\times100

=\frac{5798.85}{5860.205}\times 100

=98.95%